# ImageFolder: Autoregressive Image Generation with Folded Tokens

### v1

@@ 4.1 Experimental Settings
We test our ImageFolder tokenizer on the ImageNet 256x256 reconstruction and generation tasks. Our tokenizer training recipe is based on LlamaGen Sun et al. (2024). [Important] Due to the limited computational resources, we just report the results trained with a short training scheme, e.g., 200k iterations for the tokenizer. More results with advanced training recipes (more iterations, stronger discriminator, and strong perceptual loss) will be updated soon.
@@ 4.2 Performance analysis
GAN BigGAN Brock (2018) 6.95 224.5 0.89 0.38 112M -
Diff. ADM Dhariwal & Nichol (2021) 10.94 101.0 0.69 0.63 554M -
Diff. LDM-4 Rombach et al. (2022b) 3.60 247.7 - - 400M -
Diff. DiT-L/2 Peebles & Xie (2023b) 5.02 167.2 0.75 0.57 458M -
Diff. MAR-B Li et al. (2024c) 2.31 281.7 0.82 0.57 208M -
NAR MaskGIT Chang et al. (2022) 6.18 182.1 0.80 0.51 227M 256
NAR RCG (cond.) Li et al. (2024b) 3.49 215.5 - - 502M 256
AR. VQGAN Esser et al. (2021) 18.65 80.4 0.78 0.26 227M 256
AR. RQ-Transformer Lee et al. (2022b) 15.72 86.8 - - 480M 1024
AR. VAR Tian et al. (2024) 3.30 274.4 0.84 0.51 310M 680
AR. LlamaGen-L-16x16 Sun et al. (2024) 3.80 248.3 0.83 0.52 343M 256
AR. ImageFolder (ours) 3.53 277.2 0.77 0.57 362M 265
Table 4: Performance comparison on class-conditional ImageNet 256x256. [Important] Due to the computing limitation, the utilized ImageFolder tokenizer is trained with a short scheme, i.e., 200K iterations. Results with advanced training recipes as other comparable methods will be updated soon. denotes the utilized image tokenizer is trained on OpenImage Kuznetsova et al. (2020).

### v2

@@ 4.1 Experimental Settings
We test our ImageFolder tokenizer on the ImageNet 256x256 reconstruction and generation tasks. Our tokenizer training recipe is based on LlamaGen Sun et al. (2024). [Important] Due to the limited computational resources, we just report the results trained with a short training scheme, e.g., 250k iterations for the tokenizer. More results with advanced training recipes (more iterations, stronger discriminator, and strong perceptual loss) will be updated soon.
@@ 4.2 Performance analysis
Diff. ADM Dhariwal & Nichol (2021) - - 10.94 101.0 0.69 0.63 554M - 1000
Diff. LDM-4 Rombach et al. (2022b) 0.74 - 3.60 247.7 - - 400M - 250
Diff. DiT-L/2 Peebles & Xie (2023b) - - 5.02 167.2 0.75 0.57 458M - 250
Diff. MAR-B Li et al. (2024c) 1.22 - 2.31 281.7 0.82 0.57 208M - 64
NAR MaskGIT Chang et al. (2022) 2.28 - 6.18 182.1 0.80 0.51 227M 256 8
NAR RCG (cond.) Li et al. (2024b) - - 3.49 215.5 - - 502M 256 250
NAR TiTok-S-128 Yu et al. (2024d) 1.52 50.5 1.94 215.5 - - 502M 128 64
AR. VQGAN Esser et al. (2021) 7.94 - 18.65 80.4 0.78 0.26 227M 256 256
AR. RQ-Transformer Lee et al. (2022b) 1.83 - 15.72 86.8 - - 480M 1024 64
AR. LlamaGen-L Sun et al. (2024) 2.19 2.6 3.80 248.3 0.83 0.52 343M 256 256
AR. VAR Tian et al. (2024) 0.90 11.3 3.30 274.4 0.84 0.51 310M 680 10
AR. ImageFolder (ours) 0.80 58.0 2.60 295.0 0.75 0.63 362M 265 10
Table 4: Performance comparison on class-conditional ImageNet 256x256. [Important] Due to the computing limitation, the utilized ImageFolder tokenizer is trained with a short scheme, i.e., 250K iterations. Results with advanced training recipes as other comparable methods will be updated soon. denotes the utilized image tokenizer is trained on OpenImage Kuznetsova et al. (2020).